Given an arithmetic sequence, 2,−1,−4,−7,...
Firstly, Find the common difference, then, find the 120thterm, and lastly, find which term is −595.

Answer by Boreal(15235)

common difference is -3
-595 is -597 from +2 or 199 terms after it or the 200th term
an=a1+(n−1)d
a200=2+(199)(-3)=-595
-
a120=2+(119)(-3)=-355